# Guan, Basciftci, and Van Hentenryck

### v2

@@ 1 Introduction
The decisions made by the transit agencies and riders are usually studied separately. More specifically, transit agencies focus on designing networks with estimated travel demand, a subject extensively explored in literature (e.g., Borndörfer, Grötschel, and Pfetsch (2007), Stiglic et al. (2018), Mahéo, Kilby, and Van Hentenryck (2019), Cats, Gkiotsalitis, and Schöbel (2025)). However, omitting potential riders that can adopt the transit system during the network design can lead to suboptimal designs with lower rider adoption and higher costs. As an example, this phenomena is observed in designing On-Demand Multimodal Transit Systems, which is an emerging form of public transportation system that tightly integrates on-demand shuttles with fixed transit services such as high-frequency buses and urban rail systems. Specifically, the transit agency tends to have higher costs and can provide poor quality of service for certain riders if the behavior of latent riders are omitted during the network design resulting in non-effective utilization of on-demand shuttles (Basciftci and Van Hentenryck 2020, 2023). When it comes to modeling the choices of the riders, research leans towards the utilization of logit and machine learning models for demand analysis (Xie, Lu, and Parkany 2003, Hagenauer and Helbich 2017, Zhao et al. 2020). However, it remains critical to integrate rider choices to the transit network design problems while taking into account the decision making processes of both the transit agency and riders. Consequently, this paper explores the possibilities of integrating the aforementioned two pieces into one single framework through a game theoretic approach. Specifically, the paper proposes a bilevel optimization framework which is called Transit Network Design with Adoptions (TN-DA), with a rider choice modeling integrated as a black-box-like component. The optimal solution of the TN-DA problem can be viewed as the equilibrium point between the transit agency and the latent riders. Furthermore, the TN-DA framework is highly generalized and can be tailored to various network design problems.
@@ 2 Related Literature
A bilevel formulation is a mathematical program where certain variables are constrained by the solution of another optimization problem, as described by Kleinert et al. (2021). This approach is valuable for modeling hierarchical decision processes and has extensive applications in transportation. Among the studies that consider latent ridership, Basciftci and Van Hentenryck (2020, 2023) introduced rider adoption awareness to the design of On-Demand Multimodal Transit System (ODMTS). These studies indicate that a bilevel optimization framework needs to be utilized to model the problem to ensure a fair design for the transit system that provides accessible transit services to the areas with existing demand and potential demand. The optimization model includes a personalized choice model which associates adoption choices with the time and cost of trips in the ODMTS. These two studies and a follow-up study (Guan, Basciftci, and Van Hentenryck 2024) provide significant insights into the design of methods that solve the problem under different choice model assumptions. However, these methods still encounter computational challenges in efficiently solving large-scale problems. Building on these studies, this paper presents a generalized bilevel formulation applicable to a broad range of transit network design problems with latent demand by allowing more comprehensive choice and price models, while proposing heuristic algorithms providing desirable properties of the optimal designs with improved computational efficiency.
@@
In general, iterative approaches are widely applied to the field of transportation science, especially for complex large-scale problems. For example, the following studies investigated iterative approaches on problems related to large-scale network design such as designing distribution networks and planning public transit for large urban areas (Gattermann, Schiewe, and Schöbel 2016, Schöbel 2017, Cipriani, Gori, and Petrelli 2012, Liu et al. 2022). Studies presented by Siebert and Goerigk (2013) and Polinder et al. (2022) demonstrate a typical method for employing iterative algorithms in network design, which involves designing a network, establishing timetables, and then iteratively refining the design. When expanding the network design problem from city-wise to country-wise, another example of using iterative approaches to solve large-scale public transit problems is the study presented by Liu et al. (2019). Similar to ODMTS that will be investigated in Section 5, the goal of the problem is to optimize both the operational costs and passenger travel convenience. However, instead of using time to directly measure travel convenience, a conversion from travel time to operational income is utilized. Two case studies using the Chinese high-speed rail data presented by the authors demonstrate that iterative algorithms can efficiently solve large-scale public transit planning problems while providing profitable solutions. Furthermore, iterative approaches are utilized to solve more general network design problems such as the incomplete hub location problems—selecting hub locations without the restriction of fully-connected networks (Dai et al. 2019) and the fleet assignment problem—assigning a fleet while maximizing expected revenue (Dumas, Aithnard, and Soumis 2009). The latter study inspired Kroon, Maróti, and Nielsen (2015), and they proposed a three-step iterative approach to optimize the rolling stock to the disrupted passenger flows. Although these studies do not explicitly address latent demand in network design and omit investigating properties of optimal solutions, they highlight the efficiency of using iterative algorithms in designing large-scale logistic and urban transit networks.
@@
Besides designing public transit networks, iterative approaches are also applied to other topics in the field of transportation science when the original problem has a bilevel nature. Yu, Yang, and Yao (2010) first proposed a bilevel optimization model to determine the optimal bus frequencies to minimize the total passengers’ travel time, and an iterative algorithm that consists of a genetic algorithm and a label-marking method is then proposed to solve this model. Tawfik, Gendron, and Limbourg (2022) provided another example of applying iterative algorithms on bilevel problems by continuing on a previous study where a bilevel model formulates service design and pricing for freight networks (Tawfik and Limbourg 2019). Lodi et al. (2016) studied game theoretic frameworks between the transit agency and bus operators for optimizing certain contractual relationships in public transit systems, and employed iterative schemes as part of their solution procedure.
@@ 3 Transit Network Design with Adoptions Problem (TN-DA)
Table 1: TN-DA Notations. (For simplicity and consistency with others, notations such as $t_{\pi^{r}}$ is written as $t_{\pi}^{r}$.)
@@ 3.2 Bilevel Optimization Framework
An instance of the TN-DA problem is the On-demand Multimodal Transit System Design with Adoptions (ODMTS-DA) problem, which was examined in previous studies by Basciftci and Van Hentenryck (2020, 2023), Guan, Basciftci, and Van Hentenryck (2024). After introducing the heuristic algorithms in the subsequent Section 4, Section 5 will revisit the ODMTS-DA problem and explore its connection to the algorithms in details. Another example of the TN-DA problem, namely the Scooters-Connected Transit Systems Design with Adoptions (SCTS-DA) problem, will be showcased in Section 6. Both case studies concentrate on emerging technologies in the evolving landscape of public transportation.
@@ 5 TN-DA Example 1: ODMTS Design with Adoptions Problem
As depicted in Figure 3, ODMTS are operated around a number of hubs that serve as stops for buses and rail. The on-demand shuttles are primarily employed as the feeders to the hubs and represent an effective solution to the first-and-last-mile problem faced by most of the transit agencies. Several simulation studies have demonstrated that ODMTS can provide considerable cost and convenience benefits in different cities under diverse circumstances (Mahéo, Kilby, and Van Hentenryck 2019, Dalmeijer and Van Hentenryck 2020, Auad et al. 2021). In practice, ODMTS were also successfully deployed in the cities of Atlanta (Van Hentenryck et al. 2023) and Savannah (Chatham Area Transit 2024) in 2022 and 2024—2025, respectively. Using the ODMTS-DA as an example of the TN-DA, the goal of this section is to demonstrate the practical advantages and outcomes of the heuristic algorithms in a real-world transit network design problem.
@@ 5.2 Existing Methods for Solving ODMTS-DA
As an alternative approach, another method, known as P-Path, reformulates the corresponding bilevel problem into a single-level Mixed-Integer Programming (MIP) problem (Guan, Basciftci, and Van Hentenryck 2024). P-Path is capable of finding optimal solutions efficiently for medium-sized ODMTS-DA instances and within a reasonable time frame for large-scale instances. However, its computational resource requirements (e.g., memory) can be substantial because of the size of the resulting MIPs. Consequently, heuristic algorithms are beneficial, as they not only provide computational efficiency but might also offer a high-quality initial solution for P-Path.
@@ 5.4 Overview on ODMTS-DA Case Studies
Extra Large Atlanta 2426 66 2424 15,478 36,283 Yes Time-Based Exact Alg., Heuristics
@@
The second case study conducted with real travel data represents a regular workday in Atlanta, Georgia, to evaluate the performance of the heuristic algorithms on large-scale instances. This case study assumes that the agency will retain the rail system, remove the existing bus services, and integrate the rail system with newly designed ODMTS. As a result, all existing four rail lines are preserved, and all 38 rail stations are designated as ODMTS hubs (set $H$). Visualizations and data processing details are available in Appendix 12. As shown by Guan, Basciftci, and Van Hentenryck (2024), P-Path can efficiently find the optimal solution for large-scale ODMTS-DA instances within a reasonable time frame and clearly outperforms the exact algorithm in computational efficiency. Therefore, in this case study, only P-Path and the heuristic algorithms are compared. Furthermore, to showcase that ODMTS-DA can take the choice model $\mathcal{C}^{r}$ as a black-box, another choice model, namely Time-Transfer-Based, is applied. In this choice model, a path $\pi$ is evaluated based on both travel duration and the number of required transfers, which is denoted by $l_{\pi}^{r}$. For instance, a “shuttle $\rightarrow$ rail $\rightarrow$ bus” path involves 2 transfers. Riders also have a transfer tolerance ($l_{ub}^{r}$), which is typically a low number and is set to 3 for all riders in this case study. The $\alpha$ value for all latent riders is set to 1.5.
@@ 5.6 Results: Large Instance in Atlanta
Table 4 presents the results from heuristic algorithms on a large-scale instance, using the optimal solution from the P-Path model as a benchmark. For trip-based iterative algorithms, three step sizes, i.e., 1000, 2000, and 3000, were tested, and the step size associated with the best design objective is reported. The results show that while the heuristics do not reach the optimal solution, they still produce high-quality designs in a very short amount of time, whereas the optimal solution requires nearly five days of computation. Note that due to the structure of the ODMTS-DFD problem, the heuristics can utilize multi-threading techniques; here, 12 threads are employed. Additionally, using the solution from Algorithm arc-S2 as a warm-start for P-Path (because it provides the best solution in a shorter amount of time) can help reduce overall computation time. However, as pointed out by Guan, Basciftci, and Van Hentenryck (2024), starting P-Path with a heuristic solution does not necessarily guarantee a faster path to the optimal solution.
One may wonder why heuristics are necessary if an optimal solution is available for instances up to a certain size, even though it is time-consuming to obtain. One of the reasons is that a transit agency may need to quickly develop a transit plan for special events, and not all locations have access to powerful computing resources capable of solving a large-scale P-Path model. In these circumstances, the heuristics can be useful. Furthermore, as noted by Guan, Basciftci, and Van Hentenryck (2024), P-Path may spend a substantial amount of computing time on closing the final 1% of the optimality gap during the branch-and-bound processes. The solution reported here required 73% of its total computation time (around 85 out of the 116.48 hours) to address this last 1%, while the heuristics can achieve a better solution within just a few minutes.
@@ 6.2 Heuristic Algorithms for SCTS-DA
This section presents the heuristic settings for SCTS-DA, the formulations and the evaluation function are documented in Appendix 10.
In the case of the trip-based iterative algorithms, the value used to rank trips, denoted as $v^{r}$, is determined based on the time traveled by e-scooters since e-scooter services typically charge riders based on the amount of time they utilize the scooters. For arc-based algorithms, the expansion rules used at iteration $k$ includes:
@@ 6.3 Results: SCTS-DA in Atlanta
Similar to the large-scale ODMTS-DA case studies explained in the previous section for Atlanta, this case study employs the same dataset and parameters for buses and rails (see Appendix 12.5). For scooters, the average speed is assumed to be $16.1\text{\,}\mathrm{k}\mathrm{m}\mathrm{/}\mathrm{h}$ ($10\text{\,}\mathrm{m}\mathrm{p}\mathrm{h}$) when traveling in Atlanta. Due to the similar structure of this problem to ODMTS-DA, a modified version P-Path is derived to seek the optimal solution of SCTS-DA, and treated as a benchmark for this setting.
@@ 7 Conclusion
This work is supported by NSF Leap-HI Grant 1854684, NSF CMMI Grant 2334302, and T-SCORE Grant 69A3552047141 from the Department of Transportation.

### v3

@@ 1 Introduction
The decisions made by the transit agencies and riders are usually studied separately. More specifically, transit agencies focus on designing networks with estimated travel demand, a subject extensively explored in literature (e.g., Borndörfer et al. (2007), Stiglic et al. (2018), Mahéo et al. (2019), Cats et al. (2025)). However, omitting potential riders that can adopt the transit system during the network design can lead to suboptimal designs with lower rider adoption and higher costs. As an example, this phenomena is observed in designing On-Demand Multimodal Transit Systems, which is an emerging form of public transportation system that tightly integrates on-demand shuttles with fixed transit services such as high-frequency buses and urban rail systems. Specifically, the transit agency tends to have higher costs and can provide poor quality of service for certain riders if the behavior of latent riders are omitted during the network design resulting in non-effective utilization of on-demand shuttles (Basciftci and Van Hentenryck 2020, 2023). When it comes to modeling the choices of the riders, research leans towards the utilization of logit and machine learning models for demand analysis (Xie et al. 2003, Hagenauer and Helbich 2017, Zhao et al. 2020). However, it remains critical to integrate rider choices to the transit network design problems while taking into account the decision making processes of both the transit agency and riders. Consequently, this paper explores the possibilities of integrating the aforementioned two pieces into one single framework through a game theoretic approach. Specifically, the paper proposes a bilevel optimization framework which is called Transit Network Design with Adoptions (TN-DA), with a rider choice modeling integrated as a black-box-like component. The optimal solution of the TN-DA problem can be viewed as the equilibrium point between the transit agency and the latent riders. Furthermore, the TN-DA framework is highly generalized and can be tailored to various network design problems.
@@ 2 Related Literature
A bilevel formulation is a mathematical program where certain variables are constrained by the solution of another optimization problem, as described by Kleinert et al. (2021). This approach is valuable for modeling hierarchical decision processes and has extensive applications in transportation. Among the studies that consider latent ridership, Basciftci and Van Hentenryck (2020, 2023) introduced rider adoption awareness to the design of On-Demand Multimodal Transit System (ODMTS). These studies indicate that a bilevel optimization framework needs to be utilized to model the problem to ensure a fair design for the transit system that provides accessible transit services to the areas with existing demand and potential demand. The optimization model includes a personalized choice model which associates adoption choices with the time and cost of trips in the ODMTS. These two studies and a follow-up study (Guan et al. 2024) provide significant insights into the design of methods that solve the problem under different choice model assumptions. However, these methods still encounter computational challenges in efficiently solving large-scale problems. Building on these studies, this paper presents a generalized bilevel formulation applicable to a broad range of transit network design problems with latent demand by allowing more comprehensive choice and price models, while proposing heuristic algorithms providing desirable properties of the optimal designs with improved computational efficiency.
@@
In general, iterative approaches are widely applied to the field of transportation science, especially for complex large-scale problems. For example, the following studies investigated iterative approaches on problems related to large-scale network design such as designing distribution networks and planning public transit for large urban areas (Gattermann et al. 2016, Schöbel 2017, Cipriani et al. 2012, Liu et al. 2022). Studies presented by Siebert and Goerigk (2013) and Polinder et al. (2022) demonstrate a typical method for employing iterative algorithms in network design, which involves designing a network, establishing timetables, and then iteratively refining the design. When expanding the network design problem from city-wise to country-wise, another example of using iterative approaches to solve large-scale public transit problems is the study presented by Liu et al. (2019). Similar to ODMTS that will be investigated in Section 5, the goal of the problem is to optimize both the operational costs and passenger travel convenience. However, instead of using time to directly measure travel convenience, a conversion from travel time to operational income is utilized. Two case studies using the Chinese high-speed rail data presented by the authors demonstrate that iterative algorithms can efficiently solve large-scale public transit planning problems while providing profitable solutions. Furthermore, iterative approaches are utilized to solve more general network design problems such as the incomplete hub location problems—selecting hub locations without the restriction of fully-connected networks (Dai et al. 2019) and the fleet assignment problem—assigning a fleet while maximizing expected revenue (Dumas et al. 2009). The latter study inspired Kroon et al. (2015), and they proposed a three-step iterative approach to optimize the rolling stock to the disrupted passenger flows. Although these studies do not explicitly address latent demand in network design and omit investigating properties of optimal solutions, they highlight the efficiency of using iterative algorithms in designing large-scale logistic and urban transit networks.
@@
Besides designing public transit networks, iterative approaches are also applied to other topics in the field of transportation science when the original problem has a bilevel nature. Yu et al. (2010) first proposed a bilevel optimization model to determine the optimal bus frequencies to minimize the total passengers’ travel time, and an iterative algorithm that consists of a genetic algorithm and a label-marking method is then proposed to solve this model. Tawfik et al. (2022) provided another example of applying iterative algorithms on bilevel problems by continuing on a previous study where a bilevel model formulates service design and pricing for freight networks (Tawfik and Limbourg 2019). Lodi et al. (2016) studied game theoretic frameworks between the transit agency and bus operators for optimizing certain contractual relationships in public transit systems, and employed iterative schemes as part of their solution procedure.
@@ 3 Transit Network Design with Adoptions Problem (TN-DA)
Table 1: TN-DA Notations. (For simplicity and consistency with others, notations such as $t_{\pi^{r}}$ are written as $t_{\pi}^{r}$.)
@@ 3.2 Bilevel Optimization Framework
An instance of the TN-DA problem is the On-demand Multimodal Transit System Design with Adoptions (ODMTS-DA) problem, which was examined in previous studies by Basciftci and Van Hentenryck (2020, 2023), Guan et al. (2024). After introducing the heuristic algorithms in the subsequent Section 4, Section 5 will revisit the ODMTS-DA problem and explore its connection to the algorithms in details. Another example of the TN-DA problem, namely the Scooters-Connected Transit Systems Design with Adoptions (SCTS-DA) problem, will be showcased in Section 6. Both case studies concentrate on emerging technologies in the evolving landscape of public transportation.
@@ 5 TN-DA Example 1: ODMTS Design with Adoptions Problem
As depicted in Figure 3, ODMTS are operated around a number of hubs that serve as stops for buses and rail. The on-demand shuttles are primarily employed as the feeders to the hubs and represent an effective solution to the first-and-last-mile problem faced by most of the transit agencies. Several simulation studies have demonstrated that ODMTS can provide considerable cost and convenience benefits in different cities under diverse circumstances (Mahéo et al. 2019, Dalmeijer and Van Hentenryck 2020, Auad et al. 2021). In practice, ODMTS were also successfully deployed in the cities of Atlanta (Van Hentenryck et al. 2023) and Savannah (Chatham Area Transit 2024) in 2022 and 2024—2025, respectively. Using the ODMTS-DA as an example of the TN-DA, the goal of this section is to demonstrate the practical advantages and outcomes of the heuristic algorithms in a real-world transit network design problem.
@@ 5.2 Existing Methods for Solving ODMTS-DA
As an alternative approach, another method, known as P-Path, reformulates the corresponding bilevel problem into a single-level Mixed-Integer Programming (MIP) problem (Guan et al. 2024). P-Path is capable of finding optimal solutions efficiently for medium-sized ODMTS-DA instances and within a reasonable time frame for large-scale instances. However, its computational resource requirements (e.g., memory) can be substantial because of the size of the resulting MIPs. Consequently, heuristic algorithms are beneficial, as they not only provide computational efficiency but might also offer a high-quality initial solution for P-Path.
@@ 5.4 Overview on ODMTS-DA Case Studies
Extra Large Atlanta 2426 66 2034 15,478 36,283 Yes Time-Based Exact Alg., Heuristics
@@
The second case study conducted with real travel data represents a regular workday in Atlanta, Georgia, to evaluate the performance of the heuristic algorithms on large-scale instances. This case study assumes that the agency will retain the rail system, remove the existing bus services, and integrate the rail system with newly designed ODMTS. As a result, all existing four rail lines are preserved, and all 38 rail stations are designated as ODMTS hubs (set $H$). Visualizations and data processing details are available in Appendix 12. As shown by Guan et al. (2024), P-Path can efficiently find the optimal solution for large-scale ODMTS-DA instances within a reasonable time frame and clearly outperforms the exact algorithm in computational efficiency. Therefore, in this case study, only P-Path and the heuristic algorithms are compared. Furthermore, to showcase that ODMTS-DA can take the choice model $\mathcal{C}^{r}$ as a black-box, another choice model, namely Time-Transfer-Based, is applied. In this choice model, a path $\pi$ is evaluated based on both travel duration and the number of required transfers, which is denoted by $l_{\pi}^{r}$. For instance, a “shuttle $\rightarrow$ rail $\rightarrow$ bus” path involves 2 transfers. Riders also have a transfer tolerance ($l_{ub}^{r}$), which is typically a low number and is set to 3 for all riders in this case study. The $\alpha$ value for all latent riders is set to 1.5.
@@ 5.6 Results: Large Instance in Atlanta
Table 4 presents the results from heuristic algorithms on a large-scale instance, using the optimal solution from the P-Path model as a benchmark. For trip-based iterative algorithms, three step sizes, i.e., 1000, 2000, and 3000, were tested, and the step size associated with the best design objective is reported. The results show that while the heuristics do not reach the optimal solution, they still produce high-quality designs in a very short amount of time, whereas the optimal solution requires nearly five days of computation. Note that due to the structure of the ODMTS-DFD problem, the heuristics can utilize multi-threading techniques; here, 12 threads are employed. Additionally, using the solution from Algorithm arc-S2 as a warm-start for P-Path (because it provides the best solution in a shorter amount of time) can help reduce overall computation time. However, as pointed out by Guan et al. (2024), starting P-Path with a heuristic solution does not necessarily guarantee a faster path to the optimal solution.
One may wonder why heuristics are necessary if an optimal solution is available for instances up to a certain size, even though it is time-consuming to obtain. One of the reasons is that a transit agency may need to quickly develop a transit plan for special events, and not all locations have access to powerful computing resources capable of solving a large-scale P-Path model. In these circumstances, the heuristics can be useful. Furthermore, as noted by Guan et al. (2024), P-Path may spend a substantial amount of computing time on closing the final 1% of the optimality gap during the branch-and-bound processes. The solution reported here required 73% of its total computation time (around 85 out of the 116.48 hours) to address this last 1%, while the heuristics can achieve a better solution within just a few minutes.
@@ 6.2 Heuristic Algorithms for SCTS-DA
This section presents the heuristic settings for SCTS-DA; the formulations and the evaluation function are documented in Appendix 10.
In the case of the trip-based iterative algorithms, the value used to rank trips, denoted as $v^{r}$, is determined based on the time traveled by e-scooters since e-scooter services typically charge riders based on the amount of time they utilize the scooters. For arc-based algorithms, the expansion rules used at iteration $k$ include:
@@ 6.3 Results: SCTS-DA in Atlanta
Similar to the large-scale ODMTS-DA case studies explained in the previous section for Atlanta, this case study employs the same dataset and parameters for buses and rails (see Appendix 12.5). For scooters, the average speed is assumed to be $16.1\text{\,}\mathrm{k}\mathrm{m}\mathrm{/}\mathrm{h}$ ($10\text{\,}\mathrm{m}\mathrm{p}\mathrm{h}$) when traveling in Atlanta, and the scooter distance bound $d_{ub}^{r}$ is set at $24\text{\,}\mathrm{k}\mathrm{m}$ ($15\text{\,}\mathrm{m}\mathrm{i}\mathrm{l}\mathrm{e}\mathrm{s}$) for all trips. Moreover, this case study also employs the congested matrices for cars and buses. Due to the similar structure of this problem to ODMTS-DA, a modified version P-Path is derived to seek the optimal solution of SCTS-DA, and treated as a benchmark for this setting.
@@ 7 Conclusion
This work was partially supported by NSF Leap-HI Grant No. 1854684, NSF Grant No. 2112533, NSF CMMI Grant No. 2434302, and T-SCORE Grant 69A3552047141 from the Department of Transportation.